# Non-singular extensions of circle-valued Morse functions

### v2

@@ 2. Preliminaries
Let $M$ be a compact manifold and $f\colon M\to S^{1}$ be a smooth map. A relation $\sim$” on $M$ is defined as follows: for $x,y\in M$, $x\sim y$ if $f(x)=f(y)$ and $x,y$ are in the same connected component of $f^{-1}(f(x))$. This is an equivalence relation on $M$ and the quotient space $W_{f}\coloneqq M/\sim$ is called the Reeb graph of $f$. Furthermore, it is easy to prove that there is a unique continuous map $\overline{f}$ that makes the following diagram commutative:
@@
We define the vertices of the Reeb graph of a circle-valued Morse function on a closed manifold as corresponding to the level sets that contain its critical points. According to [9, Proposition 3.1], the Reeb graph of such a function is a finite disjoint union of finite graphs and circles. Since we also allow submersions as circle-valued Morse functions, as noted in section 2, Reeb graphs may have circle components. Furthermore, let us define the vertices of the Reeb graph of a submersion on a manifold with non-empty boundaries, where the submersion is a circle-valued Morse function on the boundaries. In this case, we take vertices corresponding to the level sets that contain the critical points of the function on the boundaries. The Reeb graph is also a finite disjoint union of finite graphs and circles, as follows from Lemma 2.4.
@@
[…]
@@
Let $M$ be a closed orientable surface and $g\colon M\times[0,1)\to S^{1}$ be a submersion such that $f\coloneq g|_{M\times\{0\}}$ is a circle-valued Morse function. Let $p\in M\times\{0\}$ be a critical point of $f$ and $w\in T_{p}(M\times[0,1))$ be an outward vector. Note that $dg_{p}(w)$ is not zero since $g$ is a submersion. If $dg_{p}(w)$ fits the orientation of $S^{1}$, then we label the vertex $q_{f}(p)$ of the Reeb graph $W_{f}$ of $f$ by “$+$”, otherwise “--”. Note that the labels do not depend on the choice of the outward vectors.
@@
The Reeb graph of $f$, in which each vertex is labeled by “$+$” or “--” according to the above rules, is called a labeled Reeb graph of $f$ and is denoted by $W_{f}^{\pm}$.
@@
Figure 1. List of local behaviors of a collapse around the vertices of $W_{f}^{\pm}$. The signs on the symbols $M$, $N$, $S$, $G$, and $J$ correspond to the signs assigned to the vertices of $W_{f}^{\pm}$ and those symbols are adopted from [4].
@@
[…]
@@
In the context of Section 2, let $\mathscr{G}^{+}$ (resp. $\mathscr{G}^{-}$) be the subset of the vertex set of $W_{f}^{\pm}$ consisting of the vertices corresponding to $G^{+}$ (resp. $G^{-}$) as described in Figure 1. A collapse $C$ is called allowable if there exists a bijection $\gamma\colon\mathscr{G}^{+}\to\mathscr{G}^{-}$ such that $C(\gamma(a))\preceq C(a)$ for every $a\in\mathscr{G}^{+}$.
@@ 3. Examples
[…]
by the commutative diagram in Section 2, where the map $C$ is induced by the inclusion $i|_{M\times\{0\}}$. By Lemma 2.4, the Reeb graph $W_{G}$ of $G$ is a finite disjoint union of finite graphs and some circles, where the vertices correspond bijctively to those of $W_{f}^{\pm}$ and the circle components are derived from those of $W_{f}^{\pm}$. Since the local forms of $G$ around the critical points of $f$ are determined by 2.2, the behaviors of $C$ around the vertices of $W_{f}^{\pm}$ are as listed in Figure 1. Therefore, the map $C$ is a collapse.
@@
Figure 7. We depict the parts used to construct a $3$-dimensional manifold $N$. The points $p_{j}$ $(j=1,2,\ldots,6)$ lie on these parts. The cylindrical sides of $\operatorname{\rm{MOD}}(M^{-})$ and $\operatorname{\rm{MOD}}(N^{+})$ are represented by $\mu$ and $\xi$, respectively. The disk side of $\operatorname{\rm{MOD}}(N^{-})$ is denoted by $\nu$. The rectangles on $\operatorname{\rm{MOD}}(S^{+})$ and $\operatorname{\rm{MOD}}(S^{-})$ are respectively denoted by $\sigma_{j}$ and $\tau_{j}$ $(j=1,2)$. The symbols $\operatorname{\rm{MOD}}(\leavevmode\nobreak\ )$ indicate these parts as in [4].

### v3

@@ 2. Preliminaries
Let $M$ be a compact manifold and $f\colon M\to S^{1}$ be a smooth map. A relation $\sim$” on $M$ is defined as follows: for $\in M, xyx\sim y$ if $f(x)=f(y)$ and $x,y$ are in the same connected component of $f^{-1}(f(x))$. This is an equivalence relation on $M$ and the quotient space $W_{f}\coloneqq M/\sim$ is called the Reeb graph of $f$. Furthermore, it is easy to prove that there is a unique continuous map $\overline{f}$ that makes the following diagram commutative:
@@
We define the vertices of the Reeb graph of a circle-valued Morse function on a closed manifold as corresponding to the level sets that contain its critical points. According to [9, Proposition 3.1], the Reeb graph of such a function is a finite disjoint union of finite graphs and circles. Since we also allow submersions as circle-valued Morse functions, as noted in 2.2, Reeb graphs may have circle components. Furthermore, let us define the vertices of the Reeb graph of a submersion on a manifold with non-empty boundaries, where the submersion is a circle-valued Morse function on the boundaries. In this case, we take vertices corresponding to the level sets that contain the critical points of the function on the boundaries. The Reeb graph is also a finite disjoint union of finite graphs and circles, as follows from Lemma 2.4.
@@
${M\times[0,1)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ g}$$\scriptstyle{i\;}$${S^{1}}$${N\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{G}$
@@
Let $M$ be a closed orientable surface and $g\colon M\times[0,1)\to S^{1}$ be a submersion such that $f\coloneq g|_{M\times\{0\}}$ is a circle-valued Morse function. Let $p\in M\times\{0\}$ be a critical point of $f$ and $w\in T_{p}(M\times[0,1))$ be an outward vector. Note that $dg_{p}(w)$ is not zero since $g$ is a submersion. If $dg_{p}(w)$ fits the orientation of $S^{1}$, then we label the vertex $q_{f}(p)$ of the Reeb graph $W_{f}$ of $f$ by “++”, otherwise “-”. Note that the labels do not depend on the choice of the outward vectors.
@@
The Reeb graph of $f$, in which each vertex is labeled by “++” or “-” according to the above rules, is called a labeled Reeb graph of $f$ and is denoted by $W_{f}^{\pm}$.
@@
Figure 1. List of local behaviors of a collapse around the vertices of $W_{f}^{\pm}$. The signs on the symbols $M$, NN, $S$, $G$, and $J$ correspond to the signs assigned to the vertices of $W_{f}^{\pm}$ and those symbols are adopted from [4].
@@
    ${W_{f}^{\pm}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\overline{f}}$$\scriptstyle{C}$${S^{1}}$${V\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{h}$
@@
In the context of 2.8, let $\mathscr{G}^{+}$ (resp. $\mathscr{G}^{-}$) be the subset of the vertex set of $W_{f}^{\pm}$ consisting of the vertices corresponding to $G^{+}$ (resp. $G^{-}$) as described in Figure 1. A collapse $C$ is called allowable if there exists a bijection $\gamma\colon\mathscr{G}^{+}\to\mathscr{G}^{-}$ such that $C(\gamma(a))\preceq C(a)$ for every $a\in\mathscr{G}^{+}$.
@@ 3. Examples
${W_{f}^{\pm}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\overline{f}}$$\scriptstyle{C}$${S^{1}}$${W_{G}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\overline{G}}$
by the commutative diagram in 2.6, where the map $C$ is induced by the inclusion $i|_{M\times\{0\}}$. By Lemma 2.4, the Reeb graph $W_{G}$ of $G$ is a finite disjoint union of finite graphs and some circles, where the vertices correspond bijctively to those of $W_{f}^{\pm}$ and the circle components are derived from those of $W_{f}^{\pm}$. Since the local forms of $G$ around the critical points of $f$ are determined by 2.4, the behaviors of $C$ around the vertices of $W_{f}^{\pm}$ are as listed in Figure 1. Therefore, the map $C$ is a collapse.
@@
Figure 7. We depict the parts used to construct a $3$-dimensional manifold NN. The points $p_{j}$ $(j=1,2,\ldots,6)$ lie on these parts. The cylindrical sides of $\operatorname{\rm{MOD}}(M^{-})$ and $\operatorname{\rm{MOD}}(N^{+})$ are represented by $\mu$ and $\xi$, respectively. The disk side of $\operatorname{\rm{MOD}}(N^{-})$ is denoted by $\nu$. The rectangles on $\operatorname{\rm{MOD}}(S^{+})$ and $\operatorname{\rm{MOD}}(S^{-})$ are respectively denoted by $\sigma_{j}$ and $\tau_{j}$ $(j=1,2)$. The symbols $\operatorname{\rm{MOD}}(\penalty 10000\ )$ indicate these parts as in [4].